Circular No. 7194

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 F2 Julianne Dalcanton, University of Washington; S. Kent, Fermi Na-

tional Accelerator Laboratory; and S. Okamura, University of Tokyo, on behalf of the Sloan Digital Sky Survey (SDSS), report the discovery by Dalcanton of a comet on several SDSS images taken on Mar. 20 through different filters; an r′-band filter shows a tail ∼ 2′ long and a sharp nucleus inside a coma of diameter ∼ 20′′. Upon receipt at the Central Bureau yesterday of the astrometry, spanning only 72 s of time, G. V. Williams found a possible link with a single-night apparently asteroidal LINEAR object in archival data for Mar. 24; this permitted Williams to find further apparently asteroidal LINEAR observations, first on Feb. 23, then on May 12, and finally on 1998 June 18. At this point, the object was placed on the NEO Confirmation Page in expectation that additional observations would confirm the cometary nature. In response, confirming CCD observations showing cometary appearance were received from M. Tichý and Z. Moravec at Kleť on June 7.9 UT (coma diameter > 15′′, tail > 50′′ in p.a. 230^{o}) and from R. A. Koff at Thornton, CO, on June 8.2 (15′′ coma, 35′′ tail in p.a. 195^{o}). Dalcanton subsequently forwarded single-night LONEOS observations obtained on Mar. 28 and found by G. Magnier. The complete set of astrometry is given on MPEC 1999-L10.

1998/9 UT  α_{2000}  δ_{2000}  Mag. Observer
June 18.15605 12^{h}06^{m}51.^{s}99 −16^{o}53′30′′.5 17.7 LINEAR
Mar. 20.44702 15 33 02.60 + 0 13 29.7 15.5 SDSS

The following orbital elements by Williams are from 27 observations, 1998 June 18–1999 June 8 (mean residual 1′′.0):

Epoch = 1998 Aug. 15.0 TT

T = 1998 Aug. 23.9824 TT  ω = 352.^{o}3927
e = 0.998555  Ω = 210.2962 }2000.0
q = 4.718760 AU  i = 56.4215
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 1 15^{h}07.^{m}64 +11^{o}24.^{′}0 4.386 5.208 140.^{o}5 7.^{o}1 15.4
11 15 04.25 +12 10.0 4.503 5.242 132.4  8.2 15.5
21 15 01.74 +12 41.9 4.639 5.275 124.0  9.2 15.6
July 1 15 00.26 +13 00.9 4.791 5.310 115.6  9.9 15.7
11 14 59.87 +13 08.8 4.955 5.346 107.3 10.5 15.8
1999 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green